I have another one, I don't have a clue
f(x)=(x^2-6x)^2-35(X^2-6x)-200
Thank you

Answer by stanbon(75887)

f(x)=(x^2-6x)^2-35(X^2-6x)-200
------------
I suppose you are asked to solve for "x":
---
Let x^2-6x = w
---
Your problem becomes:
w^2 - 35w - 200 = 0
(w+5)(w-40) = 0
w = -5 or w = 40
---
Then solve for "x" using
x^2-6x = -5 or x^2-6x = 40
===============================
Cheers,
Stan H.
==================